OpenStudy (anonymous):

What is the equation of the line that passes through the point (6, 1) and is perpendicular to the line y = -3x + 2? Please Help!

jimthompson5910 (jim_thompson5910):

what is the slope of the given line?

OpenStudy (anonymous):

For y=-3x+t, the slope is m=-3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the same as -3/1

jimthompson5910 (jim_thompson5910):

now flip the fraction to get -1/3 then flip the sign to get +1/3 or just 1/3

jimthompson5910 (jim_thompson5910):

so the slope of ANY perpendicular line to this given line is 1/3

jimthompson5910 (jim_thompson5910):

so you now know the slope of this line, which is 1/3 and you also know that this line goes through (6,1)

jimthompson5910 (jim_thompson5910):

let me know if that's enough to get started

OpenStudy (anonymous):

Ok so, here are my answer choices just to help: y = -3x + 3 y = -3x - 9 y = 1/3 x - 1 y = 1/3 x + 19

jimthompson5910 (jim_thompson5910):

use y = mx + b

jimthompson5910 (jim_thompson5910):

m = 1/3 x = 6 y = 1

jimthompson5910 (jim_thompson5910):

y = mx + b y = (1/3)x + b ... plug in m = 1/3 1 = (1/3)(6) + b ... plug in x = 6, y = 1 now solve for b

OpenStudy (anonymous):

So it would be answer choice D?

OpenStudy (anonymous):

y = 1/3 x + 19

OpenStudy (anonymous):

Wait, I think

jimthompson5910 (jim_thompson5910):

(1/3)*6 = (1/3)*(6/1) = (1*6)/(3*1) = 6/3 = 2 so, (1/3)*6 = 2

jimthompson5910 (jim_thompson5910):

that might help

OpenStudy (anonymous):

oh, C

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

what did you get when you solved for b

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

that's not correct

jimthompson5910 (jim_thompson5910):

1 = (1/3)(6) + b 1 = 2 + b b = ???

OpenStudy (anonymous):

I suck at this I'm sorry

jimthompson5910 (jim_thompson5910):

2+b is the same as b+2

jimthompson5910 (jim_thompson5910):

1 = 2 + b 1 = b + 2 b + 2 = 1 b = ???

jimthompson5910 (jim_thompson5910):

keep practicing and you'll get better at it

OpenStudy (anonymous):

-1?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

m = 1/3 b = -1

jimthompson5910 (jim_thompson5910):

therefore y = mx + b turns into \[\large y = \frac{1}{3}x - 1\]

OpenStudy (anonymous):

Wohoo! Thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome